OpenStudy (anonymous):

M is the midpoint of. Find the coordinates of B given A(4,-2) and M(4,7).

OpenStudy (anonymous):

M is the midpoint with the line of AB with the line on top

OpenStudy (anonymous):

Do you know the formula to find the coordinate of midpoint?

OpenStudy (anonymous):

Note that the x value doesn't change, so all motion is along the y direction. The distance to the midpoint is nine, so nine more upward gives coordinate (4,16) as B.

OpenStudy (anonymous):

????

OpenStudy (anonymous):

will it be -1,2

OpenStudy (anonymous):

So you neither know the midpoint formula nor understand AnimalAin's explanation?

OpenStudy (anonymous):

From what AnimalAin explains: A(4,-2) and M(4,7) Notice both x coordinates the same => B ( 4, ...) Then since Δy = 7 + 2 = 9 => y = 7 + 9 = 16

OpenStudy (anonymous):

thus B ( 4, 16)

OpenStudy (anonymous):

there isnt a 4,16 in the multipuly choice

OpenStudy (anonymous):

Let me test by my way, but I'm pretty sure my friend's way correct!

OpenStudy (anonymous):

wil it be -1,2

OpenStudy (anonymous):

Both ways yield exactly same result B ( 4, 16)

OpenStudy (anonymous):

You must have mixed up the result from some other question :(

OpenStudy (anonymous):

yeah you right am sorry bro

OpenStudy (anonymous):

Of course, I know it happens :P And I also know myself and my friend :)